# A Survey on Knowledge Distillation of Large Language Models

### 20 Feb 2024 v1

@@ 3.1.5 Feedback
where $y$ denotes the output generated by the student model in response to $x$, and $\phi_{\text{fb}}(\cdot;\theta_{T}))$ represents providing feedback from teacher LLMs. This operation evaluates the student’s output $y$ given the input $x$, by offering assessment, corrective information, or other forms of guidance. This feedback knowledge can not only be distilled into the student to also generate feedback (such as creating a student preference model) but, more importantly, enable the student to refine its responses based on the feedback. Various methods have been explored to elicit this advanced knowledge Bai et al. (2022a); Luo et al. (2023c); Cui et al. (2023a); Kwon et al. (2023); Jiang et al. (2023b); Chen et al. (2023a); Gu et al. (2024); Agarwal et al. (2024); Chen et al. (2024a); Guo et al. (2024); Ye et al. (2023); Hong et al. (2023); Lee et al. (2023a).
Preference, as previously discussed, represents a notable form of feedback knowledge from teacher models. Various knowledge of preferences could be distilled from teachers by prompting it with specific criteria. Bai et al. (2022a) introduce RLAIF for distilling harmlessness preferences from LLMs. This involves using an SFT-trained LLM to generate response pairs for each prompt, then ranking them for harmlessness to create a preference dataset. This dataset is distilled into a Preference Model (PM), which then guides the RL training of a more harmless LLM policy. WizardMath Luo et al. (2023c) places emphasis on mathematical reasoning. They employ ChatGPT as teacher to directly provide process supervision and evaluate the correctness of each step in the generated solutions. To scale up high-quality distilled preference data, Cui et al. (2023a) develop a large-scale preference dataset for distilling better preference models, UltraFeedback. It compiles various instructions and models to produce comparative data. Then, GPT-4 is used to score candidates from various aspects of preference, including instruction-following, truthfulness, honesty and helpfulness.
@@ 3.1.6 Self-Knowledge
where $\phi_{\text{sk}}(\cdot)$ is a generalized function that represents an additional process to the self-generated outputs $y$, which could include but is not limited to filtering, rewarding, or any other mechanisms for enhancing or evaluating $y$. It could be governed by external tools or the student itself $\theta_{S}$. Recent research in this area has proposed various innovative methodologies to elicit self-knowledge, demonstrating its potential for creating more efficient and autonomous learning systems. Allen-Zhu and Li (2020); Wang et al. (2022a); Sun et al. (2024b); Yang et al. (2024a); Jung et al. (2023); Huang et al. (2023a); Gulcehre et al. (2023); Yuan et al. (2024a); Xu et al. (2023b); Zelikman et al. (2022); Chen et al. (2024b); Zheng et al. (2024); Li et al. (2024c); Zhao et al. (2024); Singh et al. (2023); Chen et al. (2024c); Hosseini et al. (2024)
@@
Note that refined self-knowledge can be iteratively acquired as the student model continuously improves, further enhancing the student’s capabilities. This is Gulcehre et al. (2023) introduces a Reinforced Self-Training (ReST) framework that cyclically alternates between Grow and Improve stages to progressively obtain better self-knowledge and refine the student model. During the Grow stage, the student model generates multiple output predictions. Then, in the Improve stage, these self-generated outputs are ranked and filtered using a scoring function. Subsequently, the language model undergoes fine-tuning on this curated dataset, employing an offline RL objective. Self-Play Chen et al. (2024b) introduces a framework resembling iterative DPO, where the language model is fine-tuned to differentiate the self-generated responses from the human-annotated data. These self-generated responses could be seen as “negative knowledge” to promote the student to better align with the target distribution. Self-Rewarding Yuan et al. (2024a) explores a novel and promising approach by utilizing the language model itself as a reward model. It employs LLM-as-a-Judge prompting to autonomously assign rewards for the self-generated responses. The entire process can then be iterated, improving instruction following and reward modeling capabilities.
@@ 3.2.1 Supervised Fine-Tuning
where $y$ is the output sequence produced by the teacher model. This simple yet highly effective technique forms the basis of numerous studies in the field. Numerous researchers have successfully employed SFT to train student models using sequences generated by teacher LLMs Taori et al. (2023); Chiang et al. (2023); Wu et al. (2023c); Xu et al. (2023a); Luo et al. (2023c). Additionally, SFT has been explored in many self-distillation works  Wang et al. (2022a); Huang et al. (2023c); Xu et al. (2023b); Zelikman et al. (2022). Due to the large number of KD works applying SFT, we only list representative ones here. More detailed works can be found in §4.
@@ 3.2.3 Reinforcement Learning
This section explores advanced methods of distilling knowledge into student models using reinforcement learning (RL). This approach is especially relevant for leveraging the feedback from teacher to train student models  Bai et al. (2022a); Cui et al. (2023a); Luo et al. (2023c); Agarwal et al. (2024); Chen et al. (2024a); Ma et al. (2023a); Pang et al. (2023); Du et al. (2023a). The RL-based distillation process typically involves two main stages:
@@
This formulation guides the reward model to correctly distinguish between more and less preferable outputs based on the teacher’s criteria. Instead of learning the instance-level rewards, RLMEC Chen et al. (2024a) adopts a different approach by training a generative reward model. It is trained on an erroneous solution rewriting data distilled from a teacher LLM. This distilled reward model can produce token-level rewards for RL training.
@@ 3.5K Wikipedia Categories +
SAIL Luo et al. (2023a) IF/RAG Alpaca Data + Web Content GPT4 LLaMA Label SFT
@@
SANDBOX Liu et al. (2023c) Value Simulation
@@
ONCE Liu et al. (2023d) Recommendation Recommendation Dataset ChatGPT LLaMA Labeling SFT
@@
WizardMath Luo et al. (2023c) Math GSM8k + MATH ChatGPT LLaMA Expansion + Feedback SFT + RL
@@
WizardCoder Luo et al. (2023b) Code Code Alpaca Data ChatGPT StarCoder Expansion SFT
@@
LLaVA Liu et al. (2023b) Vision-Language COCO GPT4 LLaMA Labeling SFT
@@ 4.1.1 Instruction Following
Instruction-following capacity enables LLMs to understand and follow user-given instructions. This ability significantly enhances human-AI interaction, allowing for seamless understanding and execution of tasks as directed by users. A primary method for acquiring this skill involves constructing instruction-like prompt-response pairs and employing Supervised Fine Tuning (SFT) for model training. Data for this purpose can be manually curated by human experts or transformed from existing NLP tasks into instructional formats with templates, such as prefacing machine translation data with ”Translate this sentence to Spanish:”. However, these approaches have limitations. Manual data creation is labor-intensive, while template-based transformation lacks diversity in instructions and may not align well with natural human input. LLMs like GPT-4 offer an efficient alternative for creating diverse and controlled SFT data by their capabilities of in-context learning and instruction following. Most relevant works use OpenAI’s GPT series models to generate prompt-response data pairs and then train the student LLMs by supervised fine-tuning Wang et al. (2022a); Taori et al. (2023); Chiang et al. (2023); Wu et al. (2023c); Xu et al. (2023a); Mukherjee et al. (2023); Mitra et al. (2023); Luo et al. (2023c); Peng et al. (2023a).
@@
Complex Instructions. Some works promote students to solve more complex instructions Xu et al. (2023a); Luo et al. (2023c, b); Guo et al. (2023c). According to Xu et al. (2023a), instruction datasets derived from human-written seeds often exhibit low to moderate complexity. To enhance the complex instruction-following capabilities of smaller models, WizardLM Xu et al. (2023a) introduces Evol-Instruct. This method gradually transforms instructions into more complex forms through a multi-step evolution process, focusing on both increasing difficulty levels and expanding the diversity of topics. They conducted four rounds of evolution using the OpenAI ChatGPT API, resulting in a dataset of 250k complex instructions. Subsequently, they trained the LLaMA 7B model, referred to as WizardLM, on this dataset. In the high-difficulty section of test instructions, WizardLM even outperformed ChatGPT, achieving a win rate 7.9% higher than ChatGPT. Zhao et al. (2023e) further conduct preliminary studies revealing the effectiveness of increasing instruction complexity. Instruction Fusion Guo et al. (2023c) further uses teacher LLMs to increase the complexity by fusing two distinct evolved instructions. Furthermore, this concept of “evolving” instructions has been extended to distill specific skills such as coding Luo et al. (2023b) and mathematics Luo et al. (2023c).
@@ 4.1.3 RAG Capbility
LLMs are known to lack the ability to utilize up-to-date knowledge, and often produce responses containing factual inaccuracies due to their sole reliance on the parametric knowledge. Retrieval-Augmented Generation (RAG) is a promising technique to decrease this issue. Handling the augmented context of retrieved information is also a non-trivial skill of LLMs. Several approaches to distill RAG capabilities have been proposed Kang et al. (2023a); Luo et al. (2023a); Asai et al. (2023).
SAIL Luo et al. (2023a) starts by retrieving search results for each training case using search APIs, creating search-augmented instructions that include both the instruction and grounding information. To encourage the language model to prioritize informative retrieval results, they input each retrieved passage along with the ground truth response into the entailment model to label each retrieval result for relevance. Subsequently, the search-augmented instructions and relevance labels are fed into teacher LLMs (like GPT-4) for generating responses. Following fine-tuning on this training set, the student model becomes proficient at denoising search results and generating accurate responses. KARD Kang et al. (2023b) distills rationales $r$ from the teacher LLM in response to questions $x$. These rationales are then utilized to train two models: a student LM and a Reranker. For training the student LM, the rationales serve as a means to retrieve relevant knowledge $d$, and the student LM is subsequently fine-tuned using the rationales alongside questions and knowledge. However, during inference, only questions are available. To address this, the Reranker is trained to mimic how the retriever scores passages with the rationale by minimizing the KL divergence between $Retriever(d|r)$ and $Reranker(d|x)$. However, the integration of a fixed number of passages in language models, without considering their necessity or relevance, can reduce versatility and lead to the generation of unhelpful responses. To equip student LMs with adaptive RAG capabilities, Self-Rag Asai et al. (2023) distills this adaptive ability from teacher LLMs into a small critic model. This critic model determines whether retrieval is necessary and evaluates the quality of the retrieved results by generating ‘reflection tokens.’ For instance, Self-Rag initiates the retrieval operation when generating the reflection token $𝑅𝑒𝑡𝑟𝑖𝑒𝑣𝑒$. To distill this critic data, GPT-4 is prompted to assess the need for retrieval using few-shot demonstrations $I$, the task input $x$, and output $y$ to predict a reflection token $r$ as follows: $p(r|I,x,y)$.
@@ 4.2.3 Value
Though slightly under-explored, aligning LLMs with human values by distilling is still possible Bai et al. (2022a); Cui et al. (2023a); Yang et al. (2024a); Sun et al. (2024b). For instance, Bai et al. (2022a) propose RLAIF, utilizing AI-generated labels to interactively improve both helpfulness and harmlessness. Sun et al. (2024b) prompt the student model with 16 principles as guidelines for generating helpful, ethical, and reliable responses. Similarly, both harmless and harmful generations could be elicited by modifying the prompts, and then are used to train the preference model Yang et al. (2024a). Cui et al. (2023a) utilize GPT-4 to rank generations regarding helpfulness, truthfulness, and honesty. Liu et al. (2023c) advance the alignment of LLMs with societal values by incorporating simulated social interactions into the training process. This approach encompasses a range of elements, including demonstrations that are both in alignment and in conflict with social norms, as well as collective ratings, in-depth feedback, and responses that are revised iteratively.
@@ 4.4.4 Recommendation
Recommender systems are integral to enhancing user experience in various online services, providing personalized content based on user preferences and behaviors. Many works have demonstrated that LLMs could be directly used as recommenders without fine-tuning Wang et al. (2023i); Dai et al. (2023c) or generate auxiliary textual features to benefit recommender systems Xi et al. (2023); Ren et al. (2023); Wei et al. (2024).  Wang et al. (2023j); Ren et al. (2023); Wei et al. (2024). However, the real-time nature of online recommender systems demands rapid response times, posing a challenge with the inherent inference latency associated with LLMs. To address this, several studies have explored ways to distill and integrate the knowledge from LLMs into recommender systems, thereby leveraging their advanced capabilities while mitigating latency issues for efficient real-time recommendations Mysore et al. (2023); Zhang et al. (2023b); Liu et al. (2023d).
Mysore et al. (2023) tackle data scarcity in narrative-driven recommendation (NDR), where users provide detailed descriptions of their preferences. They utilize GPT-3 to create synthetic narrative queries from user-item interactions via few-shot prompting, then distill this data into retrieval models for NDR. Similarly, GENRE Liu et al. (2023d) employs GPT-3.5 to augment datasets with new knowledge about news summarization, user profiles, and personalized content, aiding the training of content-based recommendation models. To bridge the gap between language models and recommender systems, some research views behavior modeling as an extension of language modeling Cui et al. (2022); Liu et al. (2023k). InstructRec Zhang et al. (2023b), for instance, interprets recommendation as instruction following. They use ChatGPT to distill a wealth of user-personalized instruction data reflecting diverse preferences and intentions based on real historical interactions. This data is then used to fine-tune a 3B student language model specifically for recommendation purposes.

### 23 Feb 2024 v2

@@ 3.1.5 Feedback
where $y$ denotes the output generated by the student model in response to $x$, and $\phi_{\text{fb}}(\cdot;\theta_{T}))$ represents providing feedback from teacher LLMs. This operation evaluates the student’s output $y$ given the input $x$, by offering assessment, corrective information, or other forms of guidance. This feedback knowledge can not only be distilled into the student to also generate feedback (such as creating a student preference model) but, more importantly, enable the student to refine its responses based on the feedback. Various methods have been explored to elicit this advanced knowledge Bai et al. (2022a); Luo et al. (2023b); Cui et al. (2023a); Kwon et al. (2023); Jiang et al. (2023b); Chen et al. (2023a); Gu et al. (2024); Agarwal et al. (2024); Chen et al. (2024b); Guo et al. (2024); Ye et al. (2023); Hong et al. (2023); Lee et al. (2023a).
Preference, as previously discussed, represents a notable form of feedback knowledge from teacher models. Various knowledge of preferences could be distilled from teachers by prompting it with specific criteria. Bai et al. (2022a) introduce RLAIF for distilling harmlessness preferences from LLMs. This involves using an SFT-trained LLM to generate response pairs for each prompt, then ranking them for harmlessness to create a preference dataset. This dataset is distilled into a Preference Model (PM), which then guides the RL training of a more harmless LLM policy. WizardMath Luo et al. (2023b) places emphasis on mathematical reasoning. They employ ChatGPT as teacher to directly provide process supervision and evaluate the correctness of each step in the generated solutions. To scale up high-quality distilled preference data, Cui et al. (2023a) develop a large-scale preference dataset for distilling better preference models, UltraFeedback. It compiles various instructions and models to produce comparative data. Then, GPT-4 is used to score candidates from various aspects of preference, including instruction-following, truthfulness, honesty and helpfulness.
@@ 3.1.6 Self-Knowledge
where $\phi_{\text{sk}}(\cdot)$ is a generalized function that represents an additional process to the self-generated outputs $y$, which could include but is not limited to filtering, rewarding, or any other mechanisms for enhancing or evaluating $y$. It could be governed by external tools or the student itself $\theta_{S}$. Recent research in this area has proposed various innovative methodologies to elicit self-knowledge, demonstrating its potential for creating more efficient and autonomous learning systems. Allen-Zhu and Li (2020); Wang et al. (2022a); Sun et al. (2024b); Yang et al. (2024a); Jung et al. (2023); Huang et al. (2023a); Gulcehre et al. (2023); Yuan et al. (2024a); Xu et al. (2023b); Zelikman et al. (2022); Chen et al. (2024a); Zheng et al. (2024); Li et al. (2024c); Zhao et al. (2024); Singh et al. (2023); Chen et al. (2024c); Hosseini et al. (2024)
@@
Note that refined self-knowledge can be iteratively acquired as the student model continuously improves, further enhancing the student’s capabilities. This is Gulcehre et al. (2023) introduces a Reinforced Self-Training (ReST) framework that cyclically alternates between Grow and Improve stages to progressively obtain better self-knowledge and refine the student model. During the Grow stage, the student model generates multiple output predictions. Then, in the Improve stage, these self-generated outputs are ranked and filtered using a scoring function. Subsequently, the language model undergoes fine-tuning on this curated dataset, employing an offline RL objective. Self-Play Chen et al. (2024a) introduces a framework resembling iterative DPO, where the language model is fine-tuned to differentiate the self-generated responses from the human-annotated data. These self-generated responses could be seen as “negative knowledge” to promote the student to better align with the target distribution. Self-Rewarding Yuan et al. (2024a) explores a novel and promising approach by utilizing the language model itself as a reward model. It employs LLM-as-a-Judge prompting to autonomously assign rewards for the self-generated responses. The entire process can then be iterated, improving instruction following and reward modeling capabilities.
@@ 3.2.1 Supervised Fine-Tuning
where $y$ is the output sequence produced by the teacher model. This simple yet highly effective technique forms the basis of numerous studies in the field. Numerous researchers have successfully employed SFT to train student models using sequences generated by teacher LLMs Taori et al. (2023); Chiang et al. (2023); Wu et al. (2023c); Xu et al. (2023a); Luo et al. (2023b). Additionally, SFT has been explored in many self-distillation works  Wang et al. (2022a); Huang et al. (2023c); Xu et al. (2023b); Zelikman et al. (2022). Due to the large number of KD works applying SFT, we only list representative ones here. More detailed works can be found in §4.
@@ 3.2.3 Reinforcement Learning
This section explores advanced methods of distilling knowledge into student models using reinforcement learning (RL). This approach is especially relevant for leveraging the feedback from teacher to train student models  Bai et al. (2022a); Cui et al. (2023a); Luo et al. (2023b); Agarwal et al. (2024); Chen et al. (2024b); Ma et al. (2023a); Pang et al. (2023); Du et al. (2023a). The RL-based distillation process typically involves two main stages:
@@
This formulation guides the reward model to correctly distinguish between more and less preferable outputs based on the teacher’s criteria. Instead of learning the instance-level rewards, RLMEC Chen et al. (2024b) adopts a different approach by training a generative reward model. It is trained on an erroneous solution rewriting data distilled from a teacher LLM. This distilled reward model can produce token-level rewards for RL training.
@@ 3.5K Wikipedia Categories +
SAIL Luo et al. (2023c) IF/RAG Alpaca Data + Web Content GPT4 LLaMA Label SFT
@@
SANDBOX Liu et al. (2023b) Value Simulation
@@
ONCE Liu et al. (2023c) Recommendation Recommendation Dataset ChatGPT LLaMA Labeling SFT
@@
WizardMath Luo et al. (2023b) Math GSM8k + MATH ChatGPT LLaMA Expansion + Feedback SFT + RL
@@
WizardCoder Luo et al. (2023a) Code Code Alpaca Data ChatGPT StarCoder Expansion SFT
@@
LLaVA Liu et al. (2023e) Vision-Language COCO GPT4 LLaMA Labeling SFT
@@ 4.1.1 Instruction Following
Instruction-following capacity enables LLMs to understand and follow user-given instructions. This ability significantly enhances human-AI interaction, allowing for seamless understanding and execution of tasks as directed by users. A primary method for acquiring this skill involves constructing instruction-like prompt-response pairs and employing Supervised Fine Tuning (SFT) for model training. Data for this purpose can be manually curated by human experts or transformed from existing NLP tasks into instructional formats with templates, such as prefacing machine translation data with ”Translate this sentence to Spanish:”. However, these approaches have limitations. Manual data creation is labor-intensive, while template-based transformation lacks diversity in instructions and may not align well with natural human input. LLMs like GPT-4 offer an efficient alternative for creating diverse and controlled SFT data by their capabilities of in-context learning and instruction following. Most relevant works use OpenAI’s GPT series models to generate prompt-response data pairs and then train the student LLMs by supervised fine-tuning Wang et al. (2022a); Taori et al. (2023); Chiang et al. (2023); Wu et al. (2023c); Xu et al. (2023a); Mukherjee et al. (2023); Mitra et al. (2023); Luo et al. (2023b); Peng et al. (2023a).
@@
Complex Instructions. Some works promote students to solve more complex instructions Xu et al. (2023a); Luo et al. (2023b, a); Guo et al. (2023c). According to Xu et al. (2023a), instruction datasets derived from human-written seeds often exhibit low to moderate complexity. To enhance the complex instruction-following capabilities of smaller models, WizardLM Xu et al. (2023a) introduces Evol-Instruct. This method gradually transforms instructions into more complex forms through a multi-step evolution process, focusing on both increasing difficulty levels and expanding the diversity of topics. They conducted four rounds of evolution using the OpenAI ChatGPT API, resulting in a dataset of 250k complex instructions. Subsequently, they trained the LLaMA 7B model, referred to as WizardLM, on this dataset. In the high-difficulty section of test instructions, WizardLM even outperformed ChatGPT, achieving a win rate 7.9% higher than ChatGPT. Zhao et al. (2023e) further conduct preliminary studies revealing the effectiveness of increasing instruction complexity. Instruction Fusion Guo et al. (2023c) further uses teacher LLMs to increase the complexity by fusing two distinct evolved instructions. Furthermore, this concept of “evolving” instructions has been extended to distill specific skills such as coding Luo et al. (2023a) and mathematics Luo et al. (2023b).
@@ 4.1.3 RAG Capbility
LLMs are known to lack the ability to utilize up-to-date knowledge, and often produce responses containing factual inaccuracies due to their sole reliance on the parametric knowledge. Retrieval-Augmented Generation (RAG) is a promising technique to decrease this issue. Handling the augmented context of retrieved information is also a non-trivial skill of LLMs. Several approaches to distill RAG capabilities have been proposed Kang et al. (2023a); Luo et al. (2023c); Asai et al. (2023).
SAIL Luo et al. (2023c) starts by retrieving search results for each training case using search APIs, creating search-augmented instructions that include both the instruction and grounding information. To encourage the language model to prioritize informative retrieval results, they input each retrieved passage along with the ground truth response into the entailment model to label each retrieval result for relevance. Subsequently, the search-augmented instructions and relevance labels are fed into teacher LLMs (like GPT-4) for generating responses. Following fine-tuning on this training set, the student model becomes proficient at denoising search results and generating accurate responses. KARD Kang et al. (2023b) distills rationales $r$ from the teacher LLM in response to questions $x$. These rationales are then utilized to train two models: a student LM and a Reranker. For training the student LM, the rationales serve as a means to retrieve relevant knowledge $d$, and the student LM is subsequently fine-tuned using the rationales alongside questions and knowledge. However, during inference, only questions are available. To address this, the Reranker is trained to mimic how the retriever scores passages with the rationale by minimizing the KL divergence between $Retriever(d|r)$ and $Reranker(d|x)$. However, the integration of a fixed number of passages in language models, without considering their necessity or relevance, can reduce versatility and lead to the generation of unhelpful responses. To equip student LMs with adaptive RAG capabilities, Self-Rag Asai et al. (2023) distills this adaptive ability from teacher LLMs into a small critic model. This critic model determines whether retrieval is necessary and evaluates the quality of the retrieved results by generating ‘reflection tokens.’ For instance, Self-Rag initiates the retrieval operation when generating the reflection token $𝑅𝑒𝑡𝑟𝑖𝑒𝑣𝑒$. To distill this critic data, GPT-4 is prompted to assess the need for retrieval using few-shot demonstrations $I$, the task input $x$, and output $y$ to predict a reflection token $r$ as follows: $p(r|I,x,y)$.
@@ 4.2.3 Value
Though slightly under-explored, aligning LLMs with human values by distilling is still possible Bai et al. (2022a); Cui et al. (2023a); Yang et al. (2024a); Sun et al. (2024b). For instance, Bai et al. (2022a) propose RLAIF, utilizing AI-generated labels to interactively improve both helpfulness and harmlessness. Sun et al. (2024b) prompt the student model with 16 principles as guidelines for generating helpful, ethical, and reliable responses. Similarly, both harmless and harmful generations could be elicited by modifying the prompts, and then are used to train the preference model Yang et al. (2024a). Cui et al. (2023a) utilize GPT-4 to rank generations regarding helpfulness, truthfulness, and honesty. Liu et al. (2023b) advance the alignment of LLMs with societal values by incorporating simulated social interactions into the training process. This approach encompasses a range of elements, including demonstrations that are both in alignment and in conflict with social norms, as well as collective ratings, in-depth feedback, and responses that are revised iteratively.
@@ 4.4.4 Recommendation
Recommender systems are integral to enhancing user experience in various online services, providing personalized content based on user preferences and behaviors. Many works have demonstrated that LLMs could be directly used as recommenders without fine-tuning Wang et al. (2023i); Dai et al. (2023c) or generate auxiliary textual features to benefit recommender systems Xi et al. (2023); Ren et al. (2023); Wei et al. (2024).  Wang et al. (2023j); Ren et al. (2023); Wei et al. (2024). However, the real-time nature of online recommender systems demands rapid response times, posing a challenge with the inherent inference latency associated with LLMs. To address this, several studies have explored ways to distill and integrate the knowledge from LLMs into recommender systems, thereby leveraging their advanced capabilities while mitigating latency issues for efficient real-time recommendations Mysore et al. (2023); Zhang et al. (2023b); Liu et al. (2023c).
Mysore et al. (2023) tackle data scarcity in narrative-driven recommendation (NDR), where users provide detailed descriptions of their preferences. They utilize GPT-3 to create synthetic narrative queries from user-item interactions via few-shot prompting, then distill this data into retrieval models for NDR. Similarly, GENRE Liu et al. (2023c) employs GPT-3.5 to augment datasets with new knowledge about news summarization, user profiles, and personalized content, aiding the training of content-based recommendation models. To bridge the gap between language models and recommender systems, some research views behavior modeling as an extension of language modeling Cui et al. (2022); Liu et al. (2023k). InstructRec Zhang et al. (2023b), for instance, interprets recommendation as instruction following. They use ChatGPT to distill a wealth of user-personalized instruction data reflecting diverse preferences and intentions based on real historical interactions. This data is then used to fine-tune a 3B student language model specifically for recommendation purposes.